OpenStudy (petiteme):

A truck covers 40.0m in 8.50s while uniformly slowing down to a final velocity of 2.80m/s. (a)find the truck's original speed. (b)find its acceleration.

OpenStudy (jfraser):

use one (or more) of the equations of kinematic motion. Find the equation that has the one missing piece, \(V_0\)

OpenStudy (amistre64):

hmmm, im thinking along the lines of 2.8 = -a(8.5)+v -40 = -a(8.5)^2/2 + v(8.5)

OpenStudy (amistre64):

prolly have some negative mixed up in that

OpenStudy (anonymous):

find the truck's original speed ? but for acceleration you can find Difference of the first V and the second V because the acceleration means that and then find the mean of them . 40 m/8.5 s = 4.7m/s 4.7/2.8=1.7

OpenStudy (petiteme):

I was checking my questions and saw that I haven't answered you guys. Sorry :( And thankyou :D